OpenStudy (anonymous):

How would you simplify & find the restrictions of (2)/(x)-(2)/(x-1)+(2)/(x-2)

OpenStudy (primeralph):

equation not clear

OpenStudy (anonymous):

\[\frac{2}{x}-\frac{2}{x-2}+\frac{2}{x-2}\] \(x\neq 0, x\neq 2\) would be the restrictions

OpenStudy (anonymous):

since \(\frac{2}{x-2}-\frac{2}{x-2}=0\) you end up only with \(\frac{2}{x}\)

OpenStudy (anonymous):

Hey can anyone help me